OpenStudy (anonymous):

Solve for x: ((2/13x)+(1/4))=((95/26x)-(1/3)). I know the solution set is 6, but I want to know how to get it.

OpenStudy (anonymous):

May I help ?

OpenStudy (anonymous):

@rrey

OpenStudy (anonymous):

Sure, if you explain it well

OpenStudy (anonymous):

Am I explain last bad ?

OpenStudy (anonymous):

Kind of, but I still got it right

OpenStudy (anonymous):

Did u get any answers ?

OpenStudy (anonymous):

For the last one or this one? For the last one I got that the denominator cannot be zero, and the solution set is 6. This one, however, I can't figure out how to even start it.

OpenStudy (anonymous):

We have to get Xs in a place and numbers in another :) Got it ?:)

OpenStudy (anonymous):

yeah. But I don't know how to do that right now

OpenStudy (anonymous):

@Hero do you think you could help?

OpenStudy (anonymous):

OK ! No problem :) Look :  Just it :)

OpenStudy (anonymous):

@rrey

hero (hero):

Well, I'll go ahead and interpret it literally...

hero (hero):

You have: \[\frac{2}{13}x + \frac{1}{4} = \frac{95}{26}x - \frac{1}{3}\]

OpenStudy (anonymous):

\[\frac{ 2 }{13x }+\frac{ 1 }{ 4}=\frac{ 95 }{ 26x }-\frac{ 1 }{ }\]

OpenStudy (anonymous):

*over 3.

hero (hero):

If that's it, then you definitely posted incorrectly horizontally. Observe how wolframalpha interprets it: http://tr.im/4fmk8

OpenStudy (anonymous):

Sorry. Well anyway, I still don't know what to do. I know in the end, my solution set will be 6 and x can't equal 0. But I don't know how to do the work

hero (hero):

Every time a student posts fractions horizontally, I have to explain this before proceeding.

hero (hero):

Don't worry. I will show you.

hero (hero):

\[\frac{2}{13x}+\frac{1}{4}=\frac{95}{26x}-\frac{1}{3}\] Multiply \(\dfrac{2}{13}\) by \(\dfrac{2}{2}\) to get \[\frac{4}{26x}+\frac{1}{4}=\frac{95}{26x}-\frac{1}{3}\] Subtract \(\dfrac{4}{26}\) from both sides; Add \(\dfrac{1}{3}\) to both sides: \[\frac{1}{3}+\frac{1}{4}=\frac{95}{26x}-\frac{4}{26x}\]

hero (hero):

Then add \(\dfrac{1}{3}\) and \(\dfrac{1}{4}\) while combing fractions on the right side to get: \[\frac{7}{12} = \frac{91}{26x}\]

hero (hero):

From there, you can cross multiply to continue solving

hero (hero):

I would proceed in the following manner to make this as painless as possible: \[7(26x) = 91(12)\] \[\frac{26x}{12} = \frac{91}{7}\] I re-wrote it this way because 91 is divisible by 7 and the fraction 26/12 is reducable: \[\frac{13x}{6} = 13\] Next, clearly we can divide both sides by 13 to get \[\frac{x}{6} = \frac{13}{13}\] Which the right side obviously cancels to: \[\frac{x}{6} = 1\] and muliplying both sides by 6 yields: \[x = 6\]

OpenStudy (anonymous):

Cool. Got it. Except I just cross multiplied and then did 1092/182x

OpenStudy (anonymous):

But thank you very much. It helped a lot.

hero (hero):

I prefer to deal with the smallest possible numbers. One general rule of algebra is to simplify whenever possible. In other words, the rules of algebra encourage you to reduce whenever the opportunity presents itself in order to avoid frustration and mistakes. Furthermore, I avoid having to depend on a calculator to simplify my result.

OpenStudy (anonymous):

Hm. Well it's just freshman algebra and my teacher doesn't really restrict our calculating options so I figure I'll be okay. I'm not majoring in anything math-related anyway, but those are good tips. Thank you!